What is the acceleration due to gravity at a distance from the center of the earth which is equal to diameter of the earth


1
Expert's answer
2021-06-09T08:05:28-0400

The acceleration due to gravity is given as follows:


"g = G\dfrac{M}{R^2}"

where "G \approx 6.67\times 10^{-11} \dfrac{m^3}{kg\cdot s^2}" is the gravitational constant, "M \approx 5.97\times 10^{24}kg" is the mass of the Earth, "R" is the distance from the center of the Earth. Substituting "R = 2\cdot 6.37\times 10^{6}m" (diameter of the Earth), obtain:


"g = G\dfrac{M}{R^2}\\
g = 6.67\times 10^{-11}\cdot \dfrac{5.97\times 10^{24}}{(2\cdot 6.37\times 10^{6})^2} \approx 2.45\space m/s^2"

Answer. "2.45\space m/s^2".